# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v3

@@ 3. Classification of orbits
and $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)=-x_{\alpha_{3}}^{2}$, the pair $(\alpha_{2},\alpha_{2}+2\alpha_{3})$ enters $\zeta^{\prime}_{D,\alpha_{2}}$. This pair also belongs to $\zeta^{\prime}_{D,\alpha_{3}+\alpha_{4}}$ and $\zeta^{\prime}_{D,\alpha_{4}}$, but $\zeta^{\prime}_{D,\alpha_{4}}$ contains one more nontrivial pair, namely, $(\alpha_{4},\alpha_{3}+\alpha_{4})$, because $T_{\alpha_{4},\alpha_{3}+\alpha_{4}}(x)=x_{\alpha_{3}}$. The remaining pairs of the form $(\gamma,\alpha)$, where $\alpha\in D$ and $\gamma$ satisfies condition (1), are trivial, therefore, they lie in all $\zeta^{\prime}_{D,\gamma}$. The lemma below highlights why we consider exactly these pairs.
Lemma 3.6. Let $\lambda\in S,D=\textup{supp}(\lambda)$, $\gamma\in\Phi^{+}$ satisfy condition (1). Let for any $\alpha\in D$ the pair $(\gamma,\alpha)$ belongs to $\zeta^{\prime}_{D,\gamma}$ and formula (2) is true for all $\gamma^{\prime}\succ\gamma$ satisfying condition (1). Then formula (2) is also true for $\lambda$, $D$ and $\gamma$.
@@
Let $\alpha\in D$, $\alpha\succ\gamma$. We have $(\gamma,\alpha)\in\zeta^{\prime}_{D,\gamma}$, hence $T_{\gamma,\alpha}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. It follows from the construction of $S$ that $\alpha$ satisfies condition (1). Hence, by our assumption, $\lambda(e_{\alpha})$ can be also expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, because
@@
Finally, $(\exp(\chi).\lambda)(e_{\gamma})-\lambda(e_{\gamma})$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, so condition (2) is satisfied, as required. $\square$
@@
Finally, we are now ready for the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
@@
Proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. If $D=\{\alpha\}$ then Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds by Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. In fact, in this case $(\exp(\chi).\lambda)(e_{\gamma})=\lambda(e_{\alpha})\cdot T_{\gamma,\alpha}(x)$. Furthermore, $T_{\gamma,\alpha}(x)=0$ for $\gamma\succ\alpha$, while $T_{\gamma,\alpha}(x)=1$ for $\gamma=\alpha$. Finally, for $\gamma\prec\alpha$ satisfying condition (1), we see that $T_{\gamma,\alpha}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
@@
Now, we will prove formula (2) step by step for all $\lambda\in S$ by subsequent consideration of supports from the list. (Note that most of the supports can be considered similarly to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, see below.) Recall that $\lambda=\sum\limits_{\beta\in\Phi^{+}}\lambda_{\beta}e_{\beta}^{*}\in\mathfrak{n}^{*},~\mu=\sum\limits_{\beta\in\Phi^{+}}\mu_{\beta}\otimes e_{\beta}^{*}=\exp(\chi).\lambda\in M^{*}$.
@@
  • Let the largest (with respect to the lexicographic order) root in the support be less than $\alpha_{2}+\alpha_{3}+\alpha_{4}$. In these case, Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for all supports and all $\gamma\in\Phi^{+}$ satisfying condition (1), so the prove is complete. (In fact, $T_{\gamma,\alpha}(x)$ for all $\gamma\in\Phi^{+}$ satisfying condition (1) are equal up to multiplication by a certain nonzero scalar to $x_{\alpha_{2}}$, $x_{\alpha_{3}}$, $x_{\alpha_{4}}$, 0 or 1.)
@@
  • Let the largest root in the support be equal to $\alpha_{2}+\alpha_{3}+\alpha_{4}$. There are only two such supports in the list: $\{\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{3}\}$. Due to the fact that the condition of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is satisfied for supports consisting of one root, the pair $(\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4})$ belongs to $\zeta^{\prime}_{D,\alpha_{3}}$ regardless of whether $\alpha_{3}$ lies in $D$ or not (the presence of simple roots in the support does not affect $\zeta^{\prime}_{D,\gamma}$). The conditions of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are also satisfied, hence there is noting to prove in this case.
@@
  • Next, let the largest root in the support be equal to $\alpha_{2}+2\alpha_{3}$. Here, the first non-trivial case arises, i.e., Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ can not be proved for all such supports just by subsistent applying of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. The support may also contain $\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{4}$. This happened due to the fact that in the table cells at positions $(\alpha_{3},\alpha_{2}+2\alpha_{3})$ and $(\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4})$ the same root $\alpha_{2}+\alpha_{3}$ is written. However, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}}(x)\neq T_{\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4}}(x)$. Hence, the following situation can occur: $\alpha_{4}$ satisfies condition (1), but we can not apply Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Therefore, here we have to compute the coordinates of $\mu$ explicitly:
@@
    This shows that formula (2) holds for root $\alpha_{4}$.
  • Now, let the largest root in the support be equal either to $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ or to $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$. In these cases, everything is also almost trivial due to Lemmas Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. The only non-trivial case is when the support $D=\{\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3},\alpha_{2}\}$. At first glance, it seems that $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ cannot be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}}$. In fact, this is not the case. From Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ we know that $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$ is expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}+\alpha_{3}+\alpha_{4}}$. Therefore, $x_{\alpha_{3}}$ can be found from $(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})$ and $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$, thus $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ contains only $x_{\alpha_{3}}$.
  • If the largest root in the support is equal to $\alpha_{1}$ or $\alpha_{1}+\alpha_{2}$, then the non-trivial case is when $D$ contains $\alpha_{1}+\alpha_{2},\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}$. This case follows immediately from the formulas in the Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and the case $D=\{\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
  • If the largest root in the support is equal to $\alpha_{1}+\alpha_{2}+\alpha_{3}$, then the all non-trivial cases can be reduced to the case $D=\{\alpha_{1}+\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4},\alpha_{2}\}$. Let us write down the coordinates we are interested in:
@@
    We conclude that for roots $\alpha_{2}$ and $\alpha_{3}+\alpha_{4}$ formula (2) holds.
  • Next, if the largest root in the support is equal to $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, then the most interesting supports are those with roots $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3}+\alpha_{4}$. One can easily show that
@@
    First, consider the case when $\lambda_{\alpha_{2}+2\alpha_{3}}=0$. Clearly, the root $\alpha_{2}$ satisfies formula (2). Since $\alpha_{2}+2\alpha_{3}\notin D$, $D$ may contain either $\alpha_{3}$ or $\alpha_{2}+\alpha_{3}$. But
@@
    One can see that in both cases formula (2) is satisfied. Second, if $\lambda_{\alpha_{2}+2\alpha_{3}}\neq 0$ formula (2) for $\alpha_{2}$ follows from the fact that $\lambda_{\alpha_{2}+2\alpha_{3}}$ can be found from $\mu_{\alpha_{2}+2\alpha_{3}}$, while $x_{\alpha_{3}}$ can be found from $\mu_{\alpha_{2}+\alpha_{3}}$.
  • The largest root in the support is equal to $\alpha_{1}+\alpha_{2}+2\alpha_{3}$. Here everything essentially boils down to the supports $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}$, $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}\}$.(All other cases can be reduced to this two supports arguing similarly to the above.)In the first case, we have the following equations on coordinates of $\mu$:
@@
    For the second support, we have the following equations on coordinates of $\mu$:
@@
  • The largest root in the support is equal to $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$. Despite the fact that there are a lot of such supports, they are all amenable to simple analysis. Each such a support does not contain the roots $\alpha_{1}+\alpha_{2}+2\alpha_{3}$, $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{1}+\alpha_{2}+\alpha_{3}$. Therefore, $x_{\alpha_{4}}$, $x_{\alpha_{3}}$ and $x_{\alpha_{3}+\alpha_{4}}$ can always be expressed as a rational function in $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ (in fact, as a polynomials with coefficients from $\overline{\mathbb{F}}[\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{-1}]$). Besides, $x_{\alpha_{1}}$ can be expressed as a rational function in $\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$. It follows that if $\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ enters a monomial in a coordinate of $\mu_{\alpha}$ then this monomial can be expressed as a rational function in coordinates $\mu_{\beta}$ with $\beta\succ\alpha$. Moreover, almost all other monomials can also be expressed as polynomials in $x_{\alpha_{4}},~x_{\alpha_{3}},~x_{\alpha_{3}+\alpha_{4}}$ and $x_{\alpha_{1}}$. Finally, those monomials that can not expressed in such a way satisfy formula (2) due to the analysis above. A similar analysis can also be done with supports where the largest root is $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$.
Further steps can be analyzed similarly: either Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ works or the equations for the coordinates can be easily explicitly computed. Thus, Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is proved. $\square$
Finally, let us formulate and prove the last theorem, which allows us to obtain the main result.
Theorem 3.7. Let $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$. Then $\overline{N}.\lambda\neq\overline{N}.\lambda^{\prime}$.
Proof. Assume the contrary. Let $\lambda\neq\lambda^{\prime}$ and at the same time $\overline{N}.\lambda=\overline{N}.\lambda^{\prime}$. Let $\gamma$ be the largest positive root such that $\lambda(e_{\gamma})\neq\lambda^{\prime}(e_{\gamma})$. Then $\gamma$ must be in the support of at least one of the linear forms $\lambda$ or $\lambda^{\prime}$. Since both of these forms belong to $S$, the root $\gamma$ satisfies condition (1). Let $\mu=\exp(\chi).\lambda,~\mu^{\prime}=\exp(\chi).\lambda^{\prime}$. It follows from Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that there exist a rational functions $F_{D,\gamma}$ and $F_{D^{\prime},\gamma}$ such that
@@
Note that if $\lambda(e_{\beta})=\lambda^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$, then $\mu(e_{\beta})=\mu^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu(e_{\gamma})-\lambda(e_{\gamma})=\mu^{\prime}(e_{\gamma})-\lambda^{\prime}(e_{\gamma})$. Hence, we may choose our functions $F_{D,\gamma}$ for various $D$ and $\gamma$ so that $F_{D,\gamma}=F_{D^{\prime},\gamma}$ whenever $\{\beta\in D\mid\beta\succ\gamma\}=\{\beta\in D^{\prime}\mid\beta\succ\gamma\}$. Let $\widetilde{\mu}=\exp(\widetilde{x}).\lambda$ be an arbitrary form from their common orbit, $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$. Then
@@
Thus, distinct forms in $S$ define have distinct orbits. Note that $N.\lambda\subseteq\overline{N}.\lambda$, so if $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$ then $N.\lambda\neq N.\lambda^{\prime}$.
It remains to prove that all coadjoint orbits are obtained in this way. Consider the orbits over the field $\mathbb{F}_{q}$. It is easy to calculate $|S|$. Let $D$ be a support from the list, $|D|=k$. There are $(q-1)^{k}$ linear forms with this support. If they all lie in $S$, then this support defines $(q-1)^{k}$ forms from $S$. If there is one restriction on the scalars, then the support defines $(q-1)^{k-1}$ forms from $S$, since the restriction means that one coordinate is expressed through others. For example,
@@
for the supports 30, 31, 32, see page Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Thus, one can easily calculate that

### v4

@@ 3. Classification of orbits
and $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)=-x_{\alpha_{3}}^{2}$, the pair $(\alpha_{2},\alpha_{2}+2\alpha_{3})$ belongs to $\zeta^{\prime}_{D,\alpha_{2}}$. This pair also lies in $\zeta^{\prime}_{D,\alpha_{3}+\alpha_{4}}$ and $\zeta^{\prime}_{D,\alpha_{4}}$. Moreover, $\zeta^{\prime}_{D,\alpha_{4}}$ contains one additional nontrivial pair, namely, $(\alpha_{4},\alpha_{3}+\alpha_{4})$, because $T_{\alpha_{4},\alpha_{3}+\alpha_{4}}(x)=x_{\alpha_{3}}$. The remaining pairs of the form $(\gamma,\alpha)$ with $\alpha\in D$ and $\gamma$ satisfies condition (1) are trivial and therefore belong to all $\zeta^{\prime}_{D,\gamma}$. The lemma below explains why we focus precisely on these pairs.
Lemma 3.6. Let $\lambda\in S,D=\textup{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1). Suppose that for every $\alpha\in D$, the pair $(\gamma,\alpha)$ belongs to $\zeta^{\prime}_{D,\gamma}$, and formula (2) holds for all $\gamma^{\prime}\succ\gamma$ that satisfy condition (1). Then formula (2) also holds for $\lambda$, $D$ and $\gamma$.
@@
Let $\alpha\in D$ with $\alpha\succ\gamma$. Since $(\gamma,\alpha)\in\zeta^{\prime}_{D,\gamma}$, the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. By the definition of $S$, the root $\alpha$ satisfies condition (1). Therefore, by the inductive hypothesis, $\lambda(e\alpha)$ can also be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, because
@@
Consequently, the difference $(\exp(\chi).\lambda)(e_{\gamma})-\lambda(e_{\gamma})$ is a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, which means condition (2) is satisfied. $\square$
@@
We are now ready to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
@@
Proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. If $D=\{\alpha\}$, then Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds by Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Indeed, in this case $(\exp(\chi).\lambda)(e_{\gamma})=\lambda(e_{\alpha})\cdot T_{\gamma,\alpha}(x)$. Moreover, $T_{\gamma,\alpha}(x)=0$ for $\gamma\succ\alpha$, while $T_{\gamma,\alpha}(x)=1$ for $\gamma=\alpha$. Finally, for $\gamma\prec\alpha$ satisfying condition (1), the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
@@
Now we will prove formula (2) step by step for all $\lambda\in S$ by considering the supports from the list one after another. (Note that most supports can be treated similarly to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$; see below.) Recall that $\lambda=\sum\limits_{\beta\in\Phi^{+}}\lambda_{\beta}e_{\beta}^{*}\in\mathfrak{n}^{*},~\mu=\sum\limits_{\beta\in\Phi^{+}}\mu_{\beta}\otimes e_{\beta}^{*}=\exp(\chi).\lambda\in M^{*}$.
@@
  • Suppose the largest root (with respect to the lexicographic order) in the support is smaller than $\alpha_{2}+\alpha_{3}+\alpha_{4}$. For all such supports and for every $\gamma\in\Phi^{+}$ satisfying condition (1), Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ applies, so the proof is complete. (Indeed, for all relevant $\gamma$, the polynomials $T_{\gamma,\alpha}(x)$ are, up to a non-zero scalar factor, equal to $x_{\alpha_{2}}$, $x_{\alpha_{3}}$, $x_{\alpha_{4}}$, 0, or 11.)
@@
  • Suppose the largest root in the support equals $\alpha_{2}+\alpha_{3}+\alpha_{4}$. There are only two such supports in the list: $\{\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{3}\}$. Since Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ covers the case of one-element supports, the pair $(\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4})$ belongs to $\zeta^{\prime}_{D,\alpha_{3}}$ regardless of whether $\alpha_{3}$ lies in $D$ (the presence of simple roots in the support does not affect $\zeta^{\prime}_{D,\gamma}$). The conditions of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are also satisfied, so nothing remains to prove in this case.
@@
  • Next, suppose the largest root in the support equals $\alpha_{2}+2\alpha_{3}$. Here the first nontrivial case appears, i.e., Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ cannot be proved for all supports of this type by a straightforward application of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Such a support may also contain $\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{4}$. The complication arises because the table entries at positions $(\alpha_{3},\alpha_{2}+2\alpha_{3})$ and $(\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4})$ both involve the same root $\alpha_{2}+\alpha_{3}$. However, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}}(x)\neq T_{\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4}}(x)$. Consequently, a situation may occur where $\alpha_{4}$ satisfies condition (1) but Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ cannot be applied directly. Therefore, in this case we must compute the coordinates of $\mu$ explicitly:
@@
    This shows that formula (2) holds for the root $\alpha_{4}$.
  • Now suppose the largest root in the support equals either $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ or $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$. In these cases, the statement also follows almost trivially from Lemmas Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. The only nontrivial situation occurs for the support $D=\{\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3},\alpha_{2}\}$. At first glance, it might seem that $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ cannot be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}}$. However, this is not the case. From Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ we know that $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}+\alpha_{3}+\alpha_{4}}$. Consequently, $x_{\alpha_{3}}$ can be recovered from $(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})$ and $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$. Since $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ depends only on $x_{\alpha_{3}}$.
  • If the largest root in the support equals $\alpha_{1}$ or $\alpha_{1}+\alpha_{2}$, then the nontrivial case is when $D$ contains $\alpha_{1}+\alpha_{2},\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}$. This case follows immediately from the formulas in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ together with the analysis of the case $D=\{\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
  • If the largest root in the support equals $\alpha_{1}+\alpha_{2}+\alpha_{3}$, then all nontrivial cases reduce to $D=\{\alpha_{1}+\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4},\alpha_{2}\}$. Let us write down the relevant coordinates:
@@
    We conclude that for the roots $\alpha_{2}$ and $\alpha_{3}+\alpha_{4}$, formula (2) holds.
  • Next, if the largest root in the support equals $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, then the most interesting supports are those containing the roots $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3}+\alpha_{4}$. One can easily show that
@@
    First, consider the case $\lambda_{\alpha_{2}+2\alpha_{3}}=0$. Then clearly the root $\alpha_{2}$ satisfies formula (2). Since $\alpha_{2}+2\alpha_{3}\notin D$, the support $D$ may contain either $\alpha_{3}$ or $\alpha_{2}+\alpha_{3}$. But
@@
    One can see that formula (2) is satisfied in both situations. Second, if $\lambda_{\alpha_{2}+2\alpha_{3}}\neq 0$, then formula (2) for $\alpha_{2}$ follows from the fact that $\lambda_{\alpha_{2}+2\alpha_{3}}$ can be obtained from $\mu_{\alpha_{2}+2\alpha_{3}}$, while $x_{\alpha_{3}}$ can be obtained from $\mu_{\alpha_{2}+\alpha_{3}}$.
  • Suppose the largest root in the support equals $\alpha_{1}+\alpha_{2}+2\alpha_{3}$. Here the analysis essentially reduces to the two supports $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}\}$.(All other cases can be reduced to these two by arguments similar to those above.)For the first support, we have the following equations for the coordinates of $\mu$:
@@
    For the second support, the equations for the coordinates of $\mu$ are:"
@@
  • Suppose the largest root in the support equals $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$. Although there are many such supports, they are all amenable to a simple analysis. Each such support does not contain the roots $\alpha_{1}+\alpha_{2}+2\alpha_{3}$, $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, and $\alpha_{1}+\alpha_{2}+\alpha_{3}$. Consequently, $x_{\alpha_{4}}$, $x_{\alpha_{3}}$, and $x_{\alpha_{3}+\alpha_{4}}$ can always be expressed as rational function of $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ (in fact, as polynomials with coefficients from $\overline{\mathbb{F}}[\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{-1}]$). Moreover, $x_{\alpha_{1}}$ can be expressed as a rational function of $\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$. It follows that if $\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ appears in a monomial in a coordinate $\mu_{\alpha}$, then that monomial can be expressed as a rational function of the coordinates $\mu_{\beta}$ with $\beta\succ\alpha$. Furthermore, almost all other monomials can also be expressed as polynomials in $x_{\alpha_{4}},~x_{\alpha_{3}},~x_{\alpha_{3}+\alpha_{4}}$ and $x_{\alpha_{1}}$. Finally, those monomials that cannot be expressed in this way satisfy formula (2) by the analysis above. A similar analysis applies to supports whose largest root is $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$.
The remaining steps can be analyzed similarly: either Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ applies, or the equations for the coordinates can be written out explicitly in a straightforward manner. Thus, Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is proved. $\square$
Finally, we formulate and prove the last theorem, which leads to the main result.
Theorem 3.7. Let $\lambda,\lambda^{\prime}\in S$ with $\lambda\neq\lambda^{\prime}$. Then $\overline{N}.\lambda\neq\overline{N}.\lambda^{\prime}$.
Proof. Assume the contrary. Let $\lambda\neq\lambda^{\prime}$ but $\overline{N}.\lambda=\overline{N}.\lambda^{\prime}$. Let $\gamma$ be the largest positive root such that $\lambda(e_{\gamma})\neq\lambda^{\prime}(e_{\gamma})$. Then $\gamma$ belongs to the support of at least one of the linear forms $\lambda$ or $\lambda^{\prime}$. Since both forms lie in $S$, the root $\gamma$ satisfies condition (1). Set $\mu=\exp(\chi).\lambda,~\mu^{\prime}=\exp(\chi).\lambda^{\prime}$. By Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, there exist rational functions $F_{D,\gamma}$ and $F_{D^{\prime},\gamma}$ such that
@@
Note that if $\lambda(e_{\beta})=\lambda^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$, then $\mu(e_{\beta})=\mu^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu(e_{\gamma})-\lambda(e_{\gamma})=\mu^{\prime}(e_{\gamma})-\lambda^{\prime}(e_{\gamma})$. Therefore, we may choose our functions $F_{D,\gamma}$ for different $D$ and $\gamma$ in such a way that $F_{D,\gamma}=F_{D^{\prime},\gamma}$ whenever $\{\beta\in D\mid\beta\succ\gamma\}=\{\beta\in D^{\prime}\mid\beta\succ\gamma\}$. Let $\widetilde{\mu}=\exp(\widetilde{x}).\lambda$ be an arbitrary element of their common orbit, where $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$. Then
@@
Thus, distinct forms in $S$ have distinct orbits. Note that $N.\lambda\subseteq\overline{N}.\lambda$, so if $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$, then $N.\lambda\neq N.\lambda^{\prime}$.
It remains to prove that all coadjoint orbits are obtained in this way. Consider the orbits over the finite field $\mathbb{F}_{q}$. One can easily compute $|S|$. Let $D$ be a support from the list, with $|D|=k$. There are $(q-1)^{k}$ linear forms with this support. If they all lie in $S$, then this support contributes $(q-1)^{k}$ forms to $S$. If there is one restriction on the scalars, then the support yields $(q-1)^{k-1}$ forms in $S$, because the restriction expresses one coordinate in terms of the others. For example,
@@
for the supports 30, 31, 32, see page Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Thus, a straightforward calculation gives